Santana Rey is considering the purchase of equipment for Business Solutions that would allow the company to add a new product to its computer furniture line. The equipment is expected to cost $329,000 and to have a seven-year life and no salvage value. It will be depreciated on a straight-line basis. Business Solutions expects to sell 100 units of the equipment’s product each year. The expected annual income related to this equipment follows.

Sales $ 384,000
Costs
Materials, labor, and overhead (except depreciation) 192,000
Depreciation on new equipment 47,000
Selling and administrative expenses 33,500
Total costs and expenses 272,500
Pretax income 111,500
Income taxes (40%) 44,600
Net income $ 66,900

Required:
(1) Compute the payback period.

Payback Period
Choose Numerator: / Choose Denominator: = Payback Period
/ = Payback period
= 0

(2) Compute the accounting rate of return for this equipment.

Accounting Rate of Return
Choose Numerator: / Choose Denominator: = Accounting Rate of Return
/ = Accounting rate of return
0
Answer #1
Payback Period= Choose Numerator/ Choose Denominator = Payback period
Expected Cost/ Annual Cash Flow = Payback period
329000/ 113900 2.888498683
* Annual Cash Flow= $66900+47000=113900
Accounting Rate of Return= Choose Numerator/ Choose Denominator = ARR
Net Income / Average Cost of Equipment = ARR
66900/ 164500 40.67%
* Average cost of Equipment = 329000/2= $164500
Dear student kindly let me know, if you are having any doubt and please mark with positive rating.